• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Fatiha

    Log entry time 16:02:34 on March25,2005

    Entry number 142401

    keyword=Shift Summary

    Shift Crew: F. Benmokhtar, P. Ulmer, X. Jiang
    ----------

    8:30-11:00: Beam studies.
    11:15: Beam back, did a Harp Scan (H. S. run 2111, coda run 2927)
    12:10: MCC called, they want to open the slit for the compton,
    it will take them 5min. After the 5min, they did other beam
    studies for 15 min.
    3:05 pm, cavity problem, beam down.
    3:33 pm, Beam back, 40 uA.


    List of runs:
    ------------
    2926...Carbon... 186 K
    2927...BeO...... Harp scan, run 2111
    2928...Junk
    2929...Carbon... 200 K ps1=50, Ramesh changed it
    2930...Carbon... 200 K ps1=100
    2931...Carbon... 200 K
    2932...Carbon... 195 K Stopped, Beam down, cavity problem.
    2933...Carbon... Goes to the next shift...